The mean weight of 20-year-old females is 126 pounds and the standard deviation is 15.7. If a sample of 25 females is selected, find the probability that the mean of the sample will be greater than 128.3 pounds. Assume that the variable is normally distributed.



1
Expert's answer
2022-03-25T06:51:59-0400

We have a normal distribution, "\mu=126, \sigma=15.7, n=25."

Let's convert it to the standard normal distribution,

"z=\cfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}" =

"=\cfrac{128.3-126}{15.7/\sqrt{25}}=0.73,"

"P(\bar{X}>128.3)=1-P(\bar{X}<128.3)=\\
=1-P(Z<0.73)="

"=1-0.7673=0.2327." ​(from z-table)